OpenStudy (anonymous):

Explain in complete sentences how u might show that radicals should be simplified before being added or subtracted to another radical, give an example where this would apply.

jhonyy9 (jhonyy9):

\[\sqrt{32}+\sqrt{8}=\sqrt{16*2}+\sqrt{4*2}=4\sqrt{2}+2\sqrt{2}=6\sqrt{2}\]

jhonyy9 (jhonyy9):

- so if you need subtracting you need making the same